It is often argued that devaluation of domestic currency acts as a tax on imports and a subsidy on exports. Explain. What would you conclude when you combine this with the Lerner Symmetry theorem ? Illustrate in general equilibrium.
Answer #1

Solution:

The devaluation of the the domestic currency country has impacts such as

  • the  government will implement a policy to encourage export of goods and discourage sale of goods on the domestic market through direct payments, low-cost loans, tax relief for exporters, or government-financed international advertising hence these are the positive out comes of subsidy on exports. An export subsidy reduces the price paid by foreign importers, which means domestic consumers pay more than foreign consumers.
  • Also devaluation of currency gives a unfair advantage in international trade and keep their prices for capital and labor artificially low hence these leads to tax on imports .

Lerner theorem explanation:

Lerner symmetry theorem states that  Export taxes and import tariff have symmetric effects on trade now as international trade depends upon relative pricing . Hence in the general equilibrium :

Let, px is price of x good and py is price of y good for export tariffs , and p*x and p*y be the tax for imports hence according to the theorem p*x/p*y=px/py when there is absence of tariff and tax.

Now when

Case 1: when the country taxes on export then ,

p*x/p*y=px/py(1+t1)

where t1 is ad valorem tariff in general equilibrium.

According to the question due to devaluation there is subsidy on export hence t1 will reduce .

Now

Case 2: Now there is tax on imports

p*x/p*y=px/py(1+t2)

where t2 is ad valorem tariff in general equilibrium.

According to this only if t1=t2 then we can say that export taxes and import tariffs have the same effect on trade but after devaluation of domestic currency there is subsidy on export and tax on import hence t1 will not be equal to t2

hence we can say that this it is a positive outcome for a country but it is not compulsary that there will always be positive outcomes there can be negative outcomes too, it may lower productivity, since imports of capital equipment and machinery may become too expensive. also Devaluation also significantly reduces the overseas purchasing power of a nation’s citizens.

If you get the answer according to the  question please give a thumbs up thankyou!!.